# JAFAR: Jack up Any Feature at Any Resolution

### v2

@@ 4.3.3 Class Activation Maps Faithfulness
Following the approach in [featup], our method can be seamlessly integrated into explainability tools such as Class Activation Maps (CAMs). Despite recent advances, CAMs are still fundamentally limited by the low-resolution feature maps produced by standard vision encoders, which hinders their ability to localize fine-grained details. By upsampling the features, our method yields sharper and more informative explanations. To assess the quality of the resulting CAMs, we adopt standard evaluation metrics from the literature: Average Drop (A.D), Average Increase (A.I), Average Gain (A.G), Coherency (Coh.), and Complexity (Cplx.).
@@
A.D ($\downarrow$) A.I ($\uparrow$) A.G ($\uparrow$) Coh. ($\uparrow$) Cplx. ($\downarrow$) ADCC ($\uparrow$) Training-free      Bilinear 19.0 18.5 3.4 88.8 59.9 61.7 (-11.6)      Large Image (x8) 48.8 12.8 2.5 67.9 38.0 59.5 (-13.8)      Strided 19.8 15.1 3.5 85.9 54.2 65.3 (-8.0) Task-Dependent      CARAFE [carafe] 49.9 4.8 1.0 66.9 34.9 59.7 (-13.6)      SAPA [sapa] 8.5 32.7 4.1 96.5 69.6 55.4 (-17.9)      Dysample [dysample] 17.8 20.0 3.8 90.2 60.7 61.6 (-11.7)      ReSFU [resfu] 14.5 24.0 3.7 92.1 64.2 59.4 (-13.9) Task-Agnostic      FeatUp [featup] 15.3 24.0 4.3 91.6 58.2 64.3 (-9.0)      LiFT [lift] 66.9 8.7 2.3 65.2 9.4 53.0 (-20.3)      JAFAR 17.4 30.9 6.5 91.4 44.1 73.3
In particular, A.D, A.I, and A.G measure how sensitive the classifier’s output is to the most salient regions of the input—an effective CAM should highlight areas that, when masked, lead to a notable change in classification confidence. Since each of these metrics captures only a single aspect of CAM quality, we also report the ADCC score—an aggregate metric proposed in [adcc] that provides a more holistic evaluation. Additional details are provided in Sec.˜A.1. As illustrated qualitatively in Fig.˜4, JAFAR generates sharper and more semantically accurate CAMs compared to all baselines. While training-free methods don’t help to recover important regions, task-dependent approaches typically produce blurrier and less precise maps. Quantitative results in Tab.˜2 further support this, with JAFAR achieving the highest score on the aggregate ADCC metric—outperforming the second-best method by 8 points, a relative improvement of 12.5%.
@@ 4.3.4 Zero-Shot Open-Vocabulary Segmentation
We further evaluate our method on a zero-shot open-vocabulary segmentation task, following the setup from [maskclip], where class labels from the dataset serve as textual inputs and predictions are made by selecting the class with the highest similarity score (argmax). Using a CLIP-ViT-B/16 backbone, this approach is entirely training-free, as it does not require a learned probing head. Results show that JAFAR significantly outperforms all baselines, with particularly strong improvements on Pascal VOC. Despite the increased difficulty of ADE20K, which includes 150 classes, our method still achieves the highest performance in both mIoU and accuracy. We report only FeatUp among the task-agnostic baselines, as it is the second-best performing method.
Table 3: Zero-Shot Open-Vocabulary Evaluation. Using MaskCLIP [maskclip] for zero-shot open-vocabulary segmentation, JAFAR consistently improves performance, indicating strong alignment with the original features.
VOC ADE20K Cityscapes Upsampling mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Nearest 24.13 30.80 9.33 24.65 19.66 50.27 Bilinear 27.87 35.27 11.03 27.78 21.56 53.21 Large Image ($\times 2$) 23.24 32.16 8.08 24.94 21.91 52.22 FeatUp [featup] 32.27 39.78 13.03 33.28 24.76 60.11 JAFAR 35.70 44.93 13.61 33.28 25.26 61.73
@@ 4.3.5 Bird’s-Eye View Segmentation
Upsampling SimpleBeV [harley2023simple] PointBeV [pointbev] BeVFormer [bevformer] Low-Res 31.75 34.89 33.72 Bilinear 33.67 36.01 34.18 FeatUp 33.95 35.38 34.01 JAFAR 36.59 37.20 36.54
Finally, we studied the impact of our upsampler in a complex training pipeline. The task, evaluated on nuScenes [nuscenes], takes several images taken from cameras as input and consists on outputting the bird’s-eye view (BeV) segmentation map. In our setup, we used a frozen DINOv2 [dinov2] backbone and trained the rest of the architecture, namely, the upsampler, the BeV encoder, and the segmentation head. This task is particularly challenging, as the model must learn to map features from the image plane to the BeV plane. To ensure a fair comparison, we also trained the architecture without an upsampler, using lower-resolution input images $(224\times 476)$. We adopted the optimization hyperparameters from PointBeV [pointbev], adjusting the batch size to 1 and training for 100 epochs. Our results show that using an upsampler consistently improves predictions, regardless of the architecture employed—SimpleBev [harley2023simple], PointBeV [pointbev], or BevFormer [bevformer]. Notably, performance improves significantly when using JAFAR as the upsampler, with mIoU gains up to +5+5 points.
@@ 6 Acknowledgment
This work has been supported by chair VISA DEEP (ANR-20-CHIA-0022) Cluster PostGenAI@Paris (ANR-23-IACL-0007, FRANCE 2030). This work has been supported by PEPR Sharp (ANR-23-PEIA-0008, FRANCE 2030).
JAFAR: Jack up Any Feature at Any Resolution
@@ A.1 Evaluation
where $\odot$ denotes element-wise multiplication. This condition implies that the CAM produced from the masked image should be identical to the original CAM, ensuring that the explanation is stable. Following the approach in [adcc], we use the Pearson Correlation Coefficient between the original CAM and the CAM obtained after masking:
@@ A.2 Baselines
  • CARAFE [carafe]: For CARAFE, we use the CARAFEPack module from MMCV [mmcv], stacking four upsampling stages with a $\times 2$ ratio each, resulting in a final feature map upsampled by a factor of $\times 16$.
  • SAPA [sapa]: We adopt the default implementation from the official SAPA repository, stacking four SAPA upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • DySample [dysample]: We adopt the default implementation from the official Dysample repository, stacking four Dysample upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • ReSFU [resfu]: We use the default implementation from the official ReSFU repository, performing a direct upsampling to the target output resolution.
  • FeatUp [featup]: For FeatUp, we use the scalable JBU variant from the official FeatUp respository stacking 4 upsampling modules to achieve a total upsampling factor of $\times 16$. We re-train FeatUp using the provided training scripts. While the original paper trains FeatUp on the COCO dataset for 2,000 steps with a batch size of 4 and evaluates on the same dataset, we train it on the ImageNet training set for 50,000 steps ($25\times$ more) using the same batch size, ensuring a fair comparison across methods.
  • LiFT [lift]: For LiFT, we slightly adapt the official implementation by resizing the intermediate representations after the downsampling module to ensure compatibility with backbones using a patch size of 14 (i.e., a downsampling factor of 14). In the official code, LiFT performs a $\times 2$ upsampling and then relies on a bilinear interpolation to upsample the features to the target output resolution.
@@ C.2 LoftUp
We present an additional comparison with LoftUp [loftup], a recent baseline which relies on segmentation masks from SAM [sam] during training and employs a two-stage pipeline that includes a self-distillation phase. For a fair evaluation, we tested JAFAR within LoftUp’s official codebase using a DINOv2 ViT-S/14 backbone.
@@ Appendix E Performance
SimpleBeV [harley2023simple] Memory (MiB) 4539 4981 5305 5453 5305
@@ Appendix F Limitations
Although our method offers significant advantages over existing upsampling approaches, it also has certain limitations. Because our approach employs a global attention mechanism, each query attends to all keys, and the number of keys increases with the resolution of the input image processed by the foundation vision encoder. As a result, the computational and memory costs can become prohibitive when dealing with large key sets (e.g., $64\times 64$ or higher). While a large receptive field is desirable, attending to every key is not strictly necessary. Alternative attention mechanisms with more localized receptive fields [swin, hassani2025generalized, pan2023slide] may achieve orders-of-magnitude efficiency improvements and represent a promising direction for future work.

### v3

@@ 4.3.3 Class Activation Maps Faithfulness
Following the approach in [14], our method can be seamlessly integrated into explainability tools such as Class Activation Maps (CAMs). Despite recent advances, CAMs are still fundamentally limited by the low-resolution feature maps produced by standard vision encoders, which hinders their ability to localize fine-grained details. By upsampling the features, our method yields sharper and more informative explanations. To assess the quality of the resulting CAMs, we adopt standard evaluation metrics from the literature: Average Drop (A.D), Average Increase (A.I), Average Gain (A.G), Coherency (Coh.), and Complexity (Cplx.).
@@
A.D ($\downarrow$) A.I ($\uparrow$) A.G ($\uparrow$) Coh. ($\uparrow$) Cplx. ($\downarrow$) ADCC ($\uparrow$) Training-free      Bilinear 19.0 18.5 3.4 88.8 59.9 61.7 (-11.6)      Large Image (x8) 48.8 12.8 2.5 67.9 38.0 59.5 (-13.8)      Strided 19.8 15.1 3.5 85.9 54.2 65.3 (-8.0) Task-Dependent      CARAFE [45] 49.9 4.8 1.0 66.9 34.9 59.7 (-13.6)      SAPA [28] 8.5 32.7 4.1 96.5 69.6 55.4 (-17.9)      Dysample [24] 17.8 20.0 3.8 90.2 60.7 61.6 (-11.7)      ReSFU [55] 14.5 24.0 3.7 92.1 64.2 59.4 (-13.9) Task-Agnostic      FeatUp [14] 15.3 24.0 4.3 91.6 58.2 64.3 (-9.0)      LiFT [40] 66.9 8.7 2.3 65.2 9.4 53.0 (-20.3)      JAFAR 17.4 30.9 6.5 91.4 44.1 73.3
In particular, A.D, A.I, and A.G measure how sensitive the classifier’s output is to the most salient regions of the input—an effective CAM should highlight areas that, when masked, lead to a notable change in classification confidence. Since each of these metrics captures only a single aspect of CAM quality, we also report the ADCC score—an aggregate metric proposed in [35] that provides a more holistic evaluation. Additional details are provided in Sec.˜A.1. As illustrated qualitatively in Fig.˜4, JAFAR generates sharper and more semantically accurate CAMs compared to all baselines. While training-free methods don’t help to recover important regions, task-dependent approaches typically produce blurrier and less precise maps. Quantitative results in Tab.˜2 further support this, with JAFAR achieving the highest score on the aggregate ADCC metric—outperforming the second-best method by 8 points, a relative improvement of 12.5%.
@@ 4.3.4 Zero-Shot Open-Vocabulary Segmentation
We further evaluate our method on a zero-shot open-vocabulary segmentation task, following the setup from [54], where class labels from the dataset serve as textual inputs and predictions are made by selecting the class with the highest similarity score (argmax). Using a CLIP-ViT-B/16 backbone, this approach is entirely training-free, as it does not require a learned probing head. Results show that JAFAR significantly outperforms all baselines, with particularly strong improvements on Pascal VOC. Despite the increased difficulty of ADE20K, which includes 150 classes, our method still achieves the highest performance in both mIoU and accuracy. We report only FeatUp among the task-agnostic baselines, as it is the second-best performing method.
Table 3: Zero-Shot Open-Vocabulary Evaluation. Using MaskCLIP [54] for zero-shot open-vocabulary segmentation, JAFAR consistently improves performance, indicating strong alignment with the original features.
VOC ADE20K Cityscapes Upsampling mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Nearest 24.13 30.80 9.33 24.65 19.66 50.27 Bilinear 27.87 35.27 11.03 27.78 21.56 53.21 Large Image ($\times 2$) 23.24 32.16 8.08 24.94 21.91 52.22 FeatUp [14] 32.27 39.78 13.03 33.28 24.76 60.11 JAFAR 35.70 44.93 13.61 33.28 25.26 61.73
@@ 4.3.5 Bird’s-Eye View Segmentation
Upsampling SimpleBeV [15] PointBeV [4] BeVFormer [22] Low-Res 31.75 34.89 33.72 Bilinear 33.67 36.01 34.18 FeatUp 33.95 35.38 34.01 JAFAR 36.59 37.20 36.54
Finally, we studied the impact of our upsampler in a complex training pipeline. The task, evaluated on nuScenes [1], takes several images taken from cameras as input and consists on outputting the bird’s-eye view (BeV) segmentation map. In our setup, we used a frozen DINOv2 [31] backbone and trained the rest of the architecture, namely, the upsampler, the BeV encoder, and the segmentation head. This task is particularly challenging, as the model must learn to map features from the image plane to the BeV plane. To ensure a fair comparison, we also trained the architecture without an upsampler, using lower-resolution input images $(224\times 476)$. We adopted the optimization hyperparameters from PointBeV [4], adjusting the batch size to 1 and training for 100 epochs. Our results show that using an upsampler consistently improves predictions, regardless of the architecture employed—SimpleBev [15], PointBeV [4], or BevFormer [22]. Notably, performance improves significantly when using JAFAR as the upsampler, with mIoU gains up to +5+5 points.
@@ 6 Acknowledgment
This work has been supported by chair VISA DEEP (ANR-20-CHIA-0022) Cluster PostGenAI@Paris (ANR-23-IACL-0007, FRANCE 2030). This work has been supported by PEPR Sharp (ANR-23-PEIA-0008, FRANCE 2030). This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011014763R1 made by GENCI.
@@ A.1 Evaluation
where $\odot$ denotes element-wise multiplication. This condition implies that the CAM produced from the masked image should be identical to the original CAM, ensuring that the explanation is stable. Following the approach in [35], we use the Pearson Correlation Coefficient between the original CAM and the CAM obtained after masking:
@@ A.2 Baselines
  • CARAFE [45]: For CARAFE, we use the CARAFEPack module from MMCV [6], stacking four upsampling stages with a $\times 2$ ratio each, resulting in a final feature map upsampled by a factor of $\times 16$.
  • SAPA [28]: We adopt the default implementation from the official SAPA repository, stacking four SAPA upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • DySample [24]: We adopt the default implementation from the official Dysample repository, stacking four Dysample upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • ReSFU [55]: We use the default implementation from the official ReSFU repository, performing a direct upsampling to the target output resolution.
  • FeatUp [14]: For FeatUp, we use the scalable JBU variant from the official FeatUp respository stacking 4 upsampling modules to achieve a total upsampling factor of $\times 16$. We re-train FeatUp using the provided training scripts. While the original paper trains FeatUp on the COCO dataset for 2,000 steps with a batch size of 4 and evaluates on the same dataset, we train it on the ImageNet training set for 50,000 steps ($25\times$ more) using the same batch size, ensuring a fair comparison across methods.
  • LiFT [40]: For LiFT, we slightly adapt the official implementation by resizing the intermediate representations after the downsampling module to ensure compatibility with backbones using a patch size of 14 (i.e., a downsampling factor of 14). In the official code, LiFT performs a $\times 2$ upsampling and then relies on a bilinear interpolation to upsample the features to the target output resolution.
@@ C.2 LoftUp
We present an additional comparison with LoftUp [18], a recent baseline which relies on segmentation masks from SAM [21] during training and employs a two-stage pipeline that includes a self-distillation phase. For a fair evaluation, we tested JAFAR within LoftUp’s official codebase using a DINOv2 ViT-S/14 backbone.
@@ Appendix E Performance
SimpleBeV [15] Memory (MiB) 4539 4981 5305 5453 5305
@@ Appendix F Limitations
Although our method offers significant advantages over existing upsampling approaches, it also has certain limitations. Because our approach employs a global attention mechanism, each query attends to all keys, and the number of keys increases with the resolution of the input image processed by the foundation vision encoder. As a result, the computational and memory costs can become prohibitive when dealing with large key sets (e.g., $64\times 64$ or higher). While a large receptive field is desirable, attending to every key is not strictly necessary. Alternative attention mechanisms with more localized receptive fields [25, 16, 32] may achieve orders-of-magnitude efficiency improvements and represent a promising direction for future work.